# Sample Compression for Self-Certified Continual Learning

### v3

@@ Abstract
Continual learning algorithms aim to learn from a sequence of tasks, making the training distribution non-stationary. The majority of existing continual learning approaches in the literature rely on heuristics and do not provide learning guarantees. In this paper, we present a new method called Continual Pick-to-Learn (CoP2L), which is able to retain the most representative samples for each task in an efficient way. CoP2L combines the Pick-to-Learn algorithm (rooted in the sample compression theory) and the experience replay continual learning scheme. This allows us to provide non-vacuous upper bounds on the generalization loss of the learned predictors, numerically computable after each task. We empirically evaluate our approach on several standard continual learning benchmarks across Class-Incremental, Task-Incremental, and Domain-Incremental settings. Our results show that CoP2L is highly competitive across all setups, often outperforming existing baselines, and significantly mitigating catastrophic forgetting compared to vanilla experience replay in the Class-Incremental setting. It is possible to leverage the bounds provided by CoP2L in practical scenarios to certify the predictor reliability on previously learned tasks, in order to improve the trustworthiness of the continual learning algorithm.
@@ 1 Introduction
A common assumption in traditional machine learning is that the underlying data distribution does not evolve with time. In Continual Learning de2021continual ; wang2023comprehensive , the goal is to develop machine learning algorithms that are able to learn under settings where this assumption is replaced by a set-up where the model is trained on an evolving training data distribution, in such a way that samples are revealed one task at a time. However, when neural networks are trained on evolving data distributions, networks tend to forget the training from earlier tasks. This fact is known as the catastrophic forgetting McCloskey1989CatastrophicII ; forgetting_french ; goodfellow2013empirical . In order to cope with forgetting, various methodologies have been developed, such as regularization based approaches ewc_kirkpatrick ; zenke2017continual ; cl_li ; mas , architectural approaches pnn_rusu ; mallya2018piggyback ; expert_aljundi ; packnet_mallya ; l2p_wang , or rehearsal based approaches rolnick2019experience ; chaudhry2019efficient ; replay_shin ; icarl_rebuffi ; der . It has been shown on various applications that rehearsal-based approaches typically perform very competitively compared to the other methods in terms of performance and forgetting DellaLibera2023CLMASRAC ; Mai2020BatchlevelER ; chaudhry2019efficient ; aljundi2019onlinecontinuallearningmaximally . However, the selection of the replay buffer typically remains a heuristic, and precise upper bounds for the generalization error are not available.
@@
Our proposed continual learning scheme builds upon the Pick-to-Learn algorithm (P2L) of Paccagnan et al. paccagnan_pick_learn_2023 . P2L is a meta-algorithm that selects a compression set, which amounts to expressing the dataset in terms of ‘core’ examples. This meta-algorithm was developed specifically with sample compression theory in mind floyd1995sample ; campi2023compression . By finding a small subset of the data such that a predictor learned on this data achieves low error on the whole training set, Pick-to-Learn enables us to compute tight upper bounds for the generalization error of the learned predictor. This ability of a learning algorithm to simultaneously output a predictor and its risk certificate has been coined as self-certified learning in the literature freund98selfbounding ; perezortiz2021tighter ; marks2025pick2learn-gp . This certification is arguably an asset to increase the trustworthiness towards machine learning systems.
@@
In this paper, we propose Continual Pick-to-Learn (CoP2L), an algorithm that leverages sample compression theory to intelligently select the training data from earlier tasks to mitigate forgetting. We derive high-confidence upper bounds on the generalization error for each task simultaneously, estimated directly from the training set. We also empirically show that CoP2L mitigates the forgetting significantly and outperforms the standard replay baseline significantly in the small replay buffer size settings. Our contributions can be summarized as follows:
@@
(1) We propose the algorithm Continual Pick-to-Learn (CoP2L), that integrates the sample compression theory within the continual learning setup. To the best of our knowledge, we are the first to integrate the theoretical results from sample compression to continual learning.
@@
(2) CoP2L is able to provide high confidence non-trivial upper bounds for the generalization error by means of the sample compression theory. We experimentally showcase that the bounds are numerically computable, have non-trivial values and follow the general error trends observed on the test set. Therefore, they can actually be used as risk certificates on the model behavior on the learned tasks, improving the trustworthiness of the continually learnt model.
@@
(3) We experimentally show with a wide range of experiments that CoP2L significantly mitigates forgetting, while also being able to obtain better or comparable performance compared with several strong continual learning baselines in most cases.
@@ 1.1 Related Work
Continual learning approaches.  In the continual learning literature, various practical approaches have been developed to mitigate forgetting. Broadly, the approaches could be divided into three categories, which include regularization approaches ewc_kirkpatrick ; zenke2017continual ; cl_li ; mas , architecture-based approaches pnn_rusu ; mallya2018piggyback ; expert_aljundi ; packnet_mallya ; l2p_wang and rehearsal based approaches rolnick2019experience ; chaudhry2019efficient ; replay_shin ; icarl_rebuffi ; der .
@@
Continual learning generalization bounds.  In recent years, generalization bounds were proposed to study the asymptotic behavior of various continual learning schemes, for example using linear models Evron2022HowCC ; Goldfarb2022AnalysisOC ; pmlr-v202-lin23f , Neural Tangent Kernel (NTK) models trained with Orthogonal Gradient Descent AbbanaBennani2020GeneralisationGF ; Doan2020ATA ; farajtabar2019orthogonalgradientdescentcontinual , regularization-based methods yin2021optimizationgeneralizationregularizationbasedcontinual , or PackNet packnet_mallya inspired continual representation learning Li2022ProvableAE . Learning algorithms with convergence guarantees are were also proposed under data assumptions, such that all target tasks lie in a same space obtainable from a (linear or non-linear) combination of some input features cao2022lifelong .
@@
Different from all the listed works above, in this paper, we provide a generalization bound that is computed from the training set. Risk certificates on the test error are provided simultaneously for each task. Furthermore, also unlike the approaches above, our non-asymptotic bound is applicable to any neural network architecture, under minimal data assumptions, and does not have limitations on the continual learning setup that is employed.
@@
Sample compression.  The sample compression theory has been shown effective for providing tight generalization bounds shah_margin-sparsity_2005 ; marchand2003set ; marchand2002set ; marchand2005learning ; shah_sample_2007 . However, most sample compress approaches are limited to low-complexity models. A notable exception is the method Pick-to-Learn (P2L), which successfully provides guarantees for deep neural networks campi2023compression ; bazinet2024 and Gaussian Processes marks2025pick2learn-gp . In this paper, we adapt the key ideas introduced in the sample compression theory to the continual learning case, which enables us to provide an upper bound on the test error from training samples.
@@ 2 Background
Continual learning.  Let us consider a series of task distributions ${\mathcal{D}}_{1},{\mathcal{D}}_{2},\ldots,{\mathcal{D}}_{T}$ on an input-output space ${\mathcal{X}}\times{\mathcal{Y}}$, where the amount of tasks $T\in{\mathbb{N}}^{*}$ can be (countably) infinite. We are given a predictive model $f_{\theta}:{\mathcal{X}}\to{\mathcal{Y}}$ (e.g., a neural network architecture), with learnable parameters $\theta\in\Theta$. From randomly initialized parameters $\theta_{0}$, the aim of the continual learning process at step $t\in\{1,\ldots,T\}$ is to update the parameters $\theta_{t-1}$ into $\theta_{t}$ by learning from a new task sample $S_{t}\sim{\mathcal{D}}_{t}$. We denote the training set at task $t$ as $S_{t}=\{({\mathbf{x}}_{t,i},y_{t,i})\}_{i=1}^{n_{t}}$, with ${\mathbf{x}}_{t,i}\in{\mathcal{X}}$ and $y_{t,i}\in{\mathcal{Y}}_{t}\subseteq{\mathcal{Y}}$; that is, the same input space ${\mathcal{X}}$ is shared among all tasks, and the output space ${\mathcal{Y}}$ may be the union of several task specific output spaces ${\mathcal{Y}}_{t}$. Given a loss function $\ell:\Theta\times{\mathcal{X}}\times{\mathcal{Y}}\to{\mathbb{R}}^{+}$, we want the last updated predictor $f_{\theta_{T}}$ to maintain a low generalization loss on all observed tasks $t\in\{1,\ldots,T\}$:
@@
The empirical loss counterparts on observed training samples are given by
^$\widehat{\mathcal{L}}_{S_{t}}(\theta_{T})\ \coloneqq\ \frac{1}{n_{t}}\sum_{i=1}^{n_{t}}\ell(\theta_{T},{\mathbf{x}}_{t,i},y_{t,i})\,.$ (2)
The challenge of continual learning lies in the fact that the learner observes the task datasets $S_{1},S_{2},\ldots,S_{t},\ldots,S_{T}$ sequentially, and we assume that the system cannot keep all observed data in memory. Nevertheless, at task $t$, we would like the system to perform well on all tasks, including the previous ones $11,\dots,t-$ 1. However, simply updating the parameters $\theta_{t-1}$ learned from a previous tasks to optimize the loss ^$\widehat{\mathcal{L}}_{S_{t}}(\theta_{t})$ on a current task $t$ would lead to catastrophic forgetting forgetting_french . Therefore, as indicated in Section 1.1, numerous empirical methods have been developed to mitigate forgetting. A simple yet very effective strategy is to simply store a small subset of data from the earlier tasks, coined as the replay buffer, and to include them in the objective function of the experience replay strategy for learning task $t$:
@@
where the second term is the loss on the replay buffer ${\mathcal{B}}$. In our experiments, we use this method as a baseline as it’s shown to be a simple and very effective continual learning strategy DellaLibera2023CLMASRAC ; Mai2020BatchlevelER .
@@
Sample compression theory.  Given a training dataset $S=\{(\boldsymbol{x}_{i},y_{i})\}_{i=1}^{n}\in(\operatorname{\mathcal{X}}\times\operatorname{\mathcal{Y}})^{n}$, a family of learnable parameters $\Theta$ and a learning algorithm $A$ such that $A(S)\in\Theta$, sample compression theory provides generalization bounds for any predictor $f_{\theta}$, with $\theta=A(S)$, on the condition that $f_{\theta}$ can be provably represented as a function of a small subset of the training dataset $S$, called the compression set, and an additional source of information, called the message. If this is the case, we call $f_{\theta}$ a sample-compressed predictor. The compression set is denoted $S^{\operatorname{\mathbf{i}}}$ and is parameterized by a vector of indices $\operatorname{\mathbf{i}}\subset\{1,\ldots,n\}$. Each vector $\operatorname{\mathbf{i}}\in\operatorname{\mathscr{P}}(n)$, with $\operatorname{\mathscr{P}}(n)$ the set of all the possible subsets of $\{1,\ldots,n\}$, is ordered such that
$\operatorname{\mathbf{i}}=(i_{1},i_{2},\ldots,i_{\operatorname{|\mathbf{i}|}})\,,\mbox{ with }1\leq i_{1}<\ldots<i_{\operatorname{|\mathbf{i}|}}\leq n\,.$ (4)
@@
The compression set is defined such that $S^{\operatorname{\mathbf{i}}}=S^{(i_{1},\ldots,i_{\operatorname{|\mathbf{i}|}})}\!=\{(\boldsymbol{x}_{i_{1}},y_{i_{1}}),\ldots,(\boldsymbol{x}_{i_{\operatorname{|\mathbf{i}|}}},y_{i_{\operatorname{|\mathbf{i}|}}})\}\!\subseteq S.$ We define the complement vector $\operatorname{\operatorname{\mathbf{i}}^{{\rm\bf c}}}$ such that $\operatorname{\mathbf{i}}\cap\operatorname{\operatorname{\mathbf{i}}^{{\rm\bf c}}}=\emptyset$ and $\operatorname{\mathbf{i}}\cup\operatorname{\operatorname{\mathbf{i}}^{{\rm\bf c}}}=\{1,\ldots,n\}$. Thus, we have the complement set $S^{\operatorname{\operatorname{\mathbf{i}}^{{\rm\bf c}}}}=S\setminus S^{\operatorname{\mathbf{i}}}$ and $|\operatorname{\operatorname{\mathbf{i}}^{{\rm\bf c}}}|=n-\operatorname{|\mathbf{i}|}$.
@@
In addition to the compression set, a message $\mu$ is sometimes needed to compress the predictor. Although generally defined as a binary sequence, the message set can be defined by any set of countable sequences of symbols. Let $\Sigma=\{\sigma_{1},\sigma_{2},\ldots,\sigma_{N}\}$ be the alphabet used to construct the messages and $\Sigma^{*}$ be the set of all possible sequences, of length 00 to $\infty$, constructed using the alphabet $\Sigma$. For all $\operatorname{\mathbf{i}}\in\operatorname{\mathscr{P}}(n)$, we choose $\mathscr{M}(\operatorname{\mathbf{i}})$ a countable subset of $\Sigma^{*}$, which represents all the possible messages that can be chosen for $\operatorname{\mathbf{i}}$.
@@
Given learned parameters $\theta=A(S)$, to prove that the predictor $f_{\theta}$ is a sample-compressed one, we need to define two functions: a compression function and a reconstruction function. The compression function $\mathscr{C}:\boldsymbol{\cup}_{1\leq n\leq\infty}(\operatorname{\mathcal{X}}\times\operatorname{\mathcal{Y}})^{n}\to\boldsymbol{\cup}_{m\leq n}(\operatorname{\mathcal{X}}\times\operatorname{\mathcal{Y}})^{m}\times\boldsymbol{\cup}_{\operatorname{\mathbf{i}}\in\operatorname{\mathscr{P}}(n)}\mathscr{M}(\operatorname{\mathbf{i}})$ outputs a compression set and a message. Then, the reconstruction function $\operatorname{\mathscr{R}}:\boldsymbol{\cup}_{m\leq n}(\operatorname{\mathcal{X}}\times\operatorname{\mathcal{Y}})^{m}\times\boldsymbol{\cup}_{\operatorname{\mathbf{i}}\in\operatorname{\mathscr{P}}(n)}\mathscr{M}(\operatorname{\mathbf{i}})\to\Theta$ is defined such that $A(S)=\operatorname{\mathscr{R}}\quantity(S^{\operatorname{\mathbf{i}}},\mu)$. The forthcoming results rely on a probability distribution over the set of sample-compressed predictors $\overline{\Theta}\subseteq\Theta$. For any sample-compressed predictor $\operatorname{\mathscr{R}}(S^{\operatorname{\mathbf{i}}},\mu)$, this data-independent distribution is expressed as $P_{\overline{\Theta}}\quantity(\operatorname{\mathscr{R}}(S^{\operatorname{\mathbf{i}}},\mu))=P_{\operatorname{\mathscr{P}}(n)}(\operatorname{\mathbf{i}})P_{\mathscr{M}(\operatorname{\mathbf{i}})}(\mu)$, with $P_{\operatorname{\mathscr{P}}(n)}$ a probability distribution over $\operatorname{\mathscr{P}}(n)$ and $P_{\mathscr{M}(\operatorname{\mathbf{i}})}$ a probability distribution over $\mathscr{M}(\operatorname{\mathbf{i}})$. Following common practices marchand2005learning ; bazinet2024 , we consider hereunder $P_{\operatorname{\mathscr{P}}(n)}(\operatorname{\mathbf{i}})={n\choose\operatorname{|\mathbf{i}|}}^{{}_{-1}}\zeta(\operatorname{|\mathbf{i}|})$, with $\zeta(k){=}\tfrac{6}{\pi^{{}_{2}}}(k+1)^{-2}$. Also, given a vector of indices $\operatorname{\mathbf{i}}$, we consider $P_{\mathscr{M}(\operatorname{\mathbf{i}})}$ to be a uniform distribution over the messages.
@@
We now present a recent sample compression generalization guarantee bazinet2024 on which we build our theoretical framework. We use the shorthand notation ^$\operatorname{\widehat{\mathcal{L}}}_{S}^{\,{}_{\operatorname{\operatorname{\mathbf{i}}^{{\rm\bf c}}}}}(\cdot)$ to denote the empirical loss on the training sample that do not belong to the compression set: $S^{\operatorname{\operatorname{\mathbf{i}}^{{\rm\bf c}}}}=S\setminus S^{\mathbf{i}}$; this is mandatory to obtain an unbiased estimate of the loss, as the reconstructed predictor $\operatorname{\mathscr{R}}(S^{\operatorname{\mathbf{i}}},\mu)$ relies on $S^{\mathbf{i}}$.
Theorem 2.1 (bazinet2024 ).
For any distribution $\operatorname{\mathcal{D}}$ over $\operatorname{\mathcal{X}}\times\operatorname{\mathcal{Y}}$, for any family of set of messages $\{\mathscr{M}(\operatorname{\mathbf{i}})|\operatorname{\mathbf{i}}{\in}\operatorname{\mathscr{P}}(n)\}$, for any deterministic reconstruction function $\operatorname{\mathscr{R}}$, for any loss $\ell:\Theta\times\operatorname{\mathcal{X}}\times\operatorname{\mathcal{Y}}{\to}[0,1]$ and for any $\delta\in(0,1]$, with probability at least $1-\delta$ over the draw of $S\sim\operatorname{\mathcal{D}}^{n}$, we have

### v4

@@ Abstract
Continual learning algorithms aim to learn from a sequence of tasks. In order to avoid catastrophic forgetting, most existing approaches rely on heuristics and do not provide computable learning guarantees. In this paper, we introduce Continual Pick-to-Learn (CoP2L), a method grounded in sample compression theory that retains representative samples for each task in a principled and efficient way. This allows us to derive non-vacuous, numerically computable upper bounds on the generalization loss of the learned predictors after each task. We evaluate CoP2L on standard continual learning benchmarks under Class-Incremental and Task-Incremental settings, showing that it effectively mitigates catastrophic forgetting. It turns out that CoP2L is empirically competitive with baseline methods while certifying predictor reliability in continual learning with a non-vacuous bound.
@@ 1 INTRODUCTION
A common assumption in traditional machine learning is that the underlying data distribution does not evolve with time. In Continual Learning (De Lange et al., 2021; Wang et al., 2024), the goal is to develop machine learning algorithms that are able to learn under settings where this assumption is replaced by a setup where the model is trained on an evolving training data distribution, in such a way that samples are revealed one task at a time. However, neural networks trained on evolving data distributions tend to forget earlier tasks, a phenomenon known as catastrophic forgetting (McCloskey and Cohen, 1989; French, 1999; Goodfellow et al., 2014). In order to cope with forgetting, various methodologies have been developed, such as regularization-based approaches (Kirkpatrick et al., 2017; Zenke et al., 2017; Li and Hoiem, 2017; Aljundi et al., 2018), architectural-based approaches (Rusu et al., 2016; Mallya et al., 2018; Aljundi et al., 2017; Mallya and Lazebnik, 2018; Wang et al., 2022; He et al., 2025; Gao et al., 2023; Zhou et al., 2025, 2024) or rehearsal-based approaches (Rolnick et al., 2019; Chaudhry et al., 2019; Shin et al., 2017; Rebuffi et al., 2017a; Buzzega et al., 2020).
(a) Class-incremental setting
(b) Task-incremental setting
Figure 1: Numerical values of the proposed generalization bounds for continual learning over 10 tasks on CIFAR100, using ViT and ResNet50 backbones. The bounds hold simultaneously for all tasks.
We develop a novel sample compression theory for continual learning. Being agnostic to the architecture, our theoretical results are well-suited for rehearsal-based approaches, which are commonly used in the continual learning literature for practical applications (Libera et al., 2023; Mai et al., 2020; Chaudhry et al., 2019; Aljundi et al., 2019). We then propose a theoretically motivated algorithm that uses the replay buffer only when necessary to mitigate the forgetting. These new results pave the way for continual learning approaches based on theoretical guarantees.
@@
More precisely, our proposed continual learning scheme builds upon the Pick-to-Learn meta-algorithm (P2L) of Paccagnan et al. (2024). This meta-algorithm was developed specifically with sample compression theory in mind (Floyd and Warmuth, 1995; Campi and Garatti, 2023). By finding a small subset of the data such that a predictor learned on this subset achieves low error on the whole training set, P2L enables computing tight upper bounds for the generalization error of the learned predictor. This ability of a learning algorithm to simultaneously output a predictor and its risk certificate has been coined as self-certified learning (Freund, 1998; Pérez-Ortiz et al., 2021; Marks and Paccagnan, 2025). This certification is arguably an asset to increase the trustworthiness of machine learning systems.
@@
The Continual Pick-to-Learn (CoP2L) [^1] Our code is available at https://anonymous.4open.science/r/CoP2L_paper_code-0058/ algorithm that we propose leverages sample compression theory to intelligently select the training data from earlier tasks to mitigate forgetting. We derive high-confidence upper bounds on the generalization error for each task simultaneously, estimated directly from the training set. We also empirically show that CoP2L substantially mitigates forgetting and is comparable with several standard baselines. Our contributions can be summarized as follows:
@@
(1) We propose the algorithm Continual Pick-to-Learn (CoP2L), that integrates the sample compression theory within the continual learning setup. To the best of our knowledge, we are the first to integrate the theoretical results from sample compression to rehearsal-based continual learning.
@@
(2) CoP2L is able to provide high confidence non-trivial upper bounds for the generalization error by means of the sample compression theory. We experimentally showcase that the bounds are numerically computable, have non-trivial values and follow the general error trends observed on the test set. Therefore, they can actually be used as risk certificates for the model’s behaviour on the learned tasks, improving the trustworthiness of the continually learnt model.
@@
(3) We experimentally show with a wide range of experiments that CoP2L significantly mitigates forgetting, while also being able to obtain comparable performance compared with several strong continual learning baselines.
@@ 1.1 Related Work
In the continual learning literature, various practical approaches have been developed to mitigate forgetting. Broadly, the approaches could be divided into three categories: 1) regularization-based approaches (Kirkpatrick et al., 2017; Zenke et al., 2017; Li and Hoiem, 2017; Aljundi et al., 2018), 2) architecture-based approaches (Rusu et al., 2016; Mallya et al., 2018; Aljundi et al., 2017; Mallya and Lazebnik, 2018; Wang et al., 2022; He et al., 2025; Gao et al., 2023; Zhou et al., 2025, 2024) and 3) rehearsal-based approaches (Rolnick et al., 2019; Chaudhry et al., 2019; Shin et al., 2017; Rebuffi et al., 2017a; Buzzega et al., 2020).
Most rehearsal-based approaches use a replay buffer that contains a small subset of the dataset over which the model continues to train, even when learning over new tasks. The buffer helps to mitigate the forgetting of a model on previous tasks. However, the effect of the buffer depends on the quality of the data chosen in the buffer. A naive approach is to sample randomly the datapoints (Mai et al., 2020; Libera et al., 2023). Non-naive methods include methods such as iCarl (Rebuffi et al., 2017a), Dark Experience Replay (Buzzega et al., 2020), Gradient Episodic Memory (Lopez-Paz and Ranzato, 2017), Example Forgetting (Benkő, 2024), Global Pseudo-Task simulation (Liu et al., 2022) and Coreset-based methods (e.g., Hao et al., 2023; Tong et al., 2025; Borsos et al., 2020; Yoon et al., 2022).
Continual learning generalization bounds.
@@
In recent years, generalization bounds were proposed to study the asymptotic behavior of various specific models, e.g., for linear models (Evron et al., 2022; Goldfarb and Hand, 2023; Lin et al., 2023; Banayeeanzade et al., 2025); for Neural Tangent Kernel models trained with Orthogonal Gradient Descent (Bennani and Sugiyama, 2020; Doan et al., 2021; Farajtabar et al., 2020); for regularization-based methods (Yin et al., 2020); or for a PackNet-inspired (Mallya and Lazebnik, 2018) model for continual representation learning (Li et al., 2022). Learning algorithms with sample complexity guarantees were also proposed under data assumptions, such that all target tasks lie in a same space obtainable from a (linear or non-linear) combination of some input features (Cao et al., 2022). Finally, Friedman and Meir (2025) presented new PAC-Bayes bounds for the backward transfer loss of stochastic models trained without a replay buffer.
@@
In contrast with the listed works above, in this paper, we simultaneously provide risk certificates on the true risk of each task, our bound is non-asymptotic and can be computed from the training set. Moreover, our bound is applicable to any neural network architecture, under minimal data assumptions and does not have limitations on the continual learning setup that is employed.
Sample compression.
@@
The sample compression theory has been shown to be effective for providing tight generalization bounds (Marchand and Shawe-Taylor, 2002; Marchand and Sokolova, 2005; Laviolette et al., 2005; Campi and Garatti, 2023). However, most sample compression approaches are limited to low-complexity models. A notable exception is the method Pick-to-Learn (P2L), presented by Paccagnan et al. (2024), which successfully provides guarantees for deep neural networks (Bazinet et al., 2025) and Gaussian Processes (Marks and Paccagnan, 2025). In this paper, we adapt the key ideas introduced in the sample compression theory to the continual learning case, which enables us to provide an upper bound on the true risk from training samples.
@@ 2 BACKGROUND
Continual learning.
Let us consider a meta-distribution $\operatorname{\mathfrak{D}}$ and a sequence of i.i.d. (independent and identically distributed) task distributions ${\mathcal{D}}_{1},{\mathcal{D}}_{2},\ldots,{\mathcal{D}}_{T}\sim\operatorname{\mathfrak{D}}$ on an input-output space ${\mathcal{X}}\times{\mathcal{Y}}$. We are given a predictive model $f_{\theta}:{\mathcal{X}}\to{\mathcal{Y}}$ (e.g., a neural network architecture), with learnable parameters $\theta\in\Theta$. From randomly initialized parameters $\theta_{0}$, the aim of the continual learning process at step $t\in\{1,\ldots,T\}$ is to update the parameters $\theta_{t-1}$ into $\theta_{t}$ by learning from a new task sample $S_{t}\sim{\mathcal{D}}_{t}$. We denote the training set at task $t$ as $S_{t}=\{({\mathbf{x}}_{t,i},y_{t,i})\}_{i=1}^{n_{t}}$, with ${\mathbf{x}}_{t,i}\in{\mathcal{X}}$ and $y_{t,i}\in{\mathcal{Y}}_{t}\subseteq{\mathcal{Y}}$; that is, the same input space $\mathcal{X}$} is shared among all tasks, and the output space $\mathcal{Y}$} may be the union of several task specific output spaces ${\mathcal{Y}}_{t}$. Given a loss function $\ell:\Theta\times{\mathcal{X}}\times{\mathcal{Y}}\to[0,1]$, we want the last updated predictor $f_{\theta_{T}}$ to maintain a low generalization loss on all observed tasks $t\in\{1,\ldots,T\}$:
@@
The empirical loss counterpart on observed training samples is given by
$\widehat{\mathcal{L}}_{S_{t}}(\theta_{T})\ \coloneqq\ \frac{1}{n_{t}}\sum_{i=1}^{n_{t}}\ell(\theta_{T},{\mathbf{x}}_{t,i},y_{t,i})\,.$ (2)
The challenge of continual learning lies in the fact that the learner observes the task datasets $S_{1},S_{2},\ldots,S_{t},\ldots,S_{T}$ sequentially and that we assume that the system cannot keep all observed data in memory. Nevertheless, at task $t$, we would like the system to perform well on all tasks, including the previous ones $1,\dots,t-1$. However, simply updating the parameters $\theta_{t-1}$ learned from a previous tasks to optimize the loss $\widehat{\mathcal{L}}_{S_{t}}(\theta_{t})$ on a current task $t$ would lead to catastrophic forgetting (French, 1999). Therefore, as indicated in Section 1.1, numerous empirical methods have been developed to mitigate forgetting. A simple yet very effective strategy is to simply store a small subset of data from the earlier tasks, coined as the replay buffer, and to include them in the objective function of the experience replay strategy for learning task $t$:
@@
where the second term is the loss on the replay buffer $\mathcal{B}$}.
Sample compression theory.
@@
Let us consider a training dataset $S=\{(\boldsymbol{x}_{i},y_{i})\}_{i=1}^{n}\in(\operatorname{\mathcal{X}}\times\operatorname{\mathcal{Y}})^{n}$ sampled i.i.d. from an unknown distribution $\operatorname{\mathcal{D}}$, a family of learnable parameters $\Theta$ and a learning algorithm $A$ such that $A(S)\in\Theta$. In this setting, sample compression theory provides generalization bounds for any predictor $f_{\theta}$, with $\theta=A(S)$, on the condition that $f_{\theta}$ can be provably represented as a function of a small subset of the training dataset $S$, called the compression set, and an additional source of information, called the message. If this is the case, we call $f_{\theta}$ a sample-compressed predictor.
@@
The compression set is denoted $S^{\operatorname{\mathbf{i}}}$ and is parameterized by a strictly increasing sequence of indices $\operatorname{\mathbf{i}}{\,\in\,}\operatorname{\mathscr{P}}(n)$, with $\operatorname{\mathscr{P}}(n)$ the powerset of $\{1,\ldots,n\}$. The compression set is defined such that $S^{\operatorname{\mathbf{i}}}=S^{(i_{1},\ldots,i_{\operatorname{|\mathbf{i}|}})}\!=\{(\boldsymbol{x}_{i_{1}},y_{i_{1}}),\ldots,(\boldsymbol{x}_{i_{\operatorname{|\mathbf{i}|}}},y_{i_{\operatorname{|\mathbf{i}|}}})\}\!\subseteq S.$ We define the complement sequence $\operatorname{\operatorname{\mathbf{i}}^{{\rm\bf c}}}$ such that $\operatorname{\mathbf{i}}\cap\operatorname{\operatorname{\mathbf{i}}^{{\rm\bf c}}}=\emptyset$ and $\operatorname{\mathbf{i}}\cup\operatorname{\operatorname{\mathbf{i}}^{{\rm\bf c}}}=\{1,\ldots,n\}$. Thus, we have the complement set $S^{\operatorname{\operatorname{\mathbf{i}}^{{\rm\bf c}}}}=S\setminus S^{\operatorname{\mathbf{i}}}$ and $\operatorname{|\mathbf{i}^{{\rm\bf c}}|}=n-\operatorname{|\mathbf{i}|}$.
@@
In addition to the compression set, a message $\mu$ could be needed to compress the predictor. Although generally defined as a binary sequence, the message set can be any set of countable sequences of symbols. Let $\Sigma=\{\sigma_{1},\sigma_{2},\ldots,\sigma_{N}\}$ be the alphabet used to construct the messages and $\Sigma^{*}$ be the set of all possible sequences, of length 0 to $\infty$, constructed using the alphabet $\Sigma$. For all $\operatorname{\mathbf{i}}\,{\in}\,\operatorname{\mathscr{P}}(n)$, we choose $\mathscr{M}(\operatorname{\mathbf{i}})$ a countable subset of $\Sigma^{*}$, which represents all the possible messages that can be chosen for $\operatorname{\mathbf{i}}$.
@@
Given learned parameters $\theta=A(S)$, to prove that the predictor $f_{\theta}$ is a sample-compressed one, we need to define two functions: a compression function and a reconstruction function. The compression function $\mathscr{C}:\boldsymbol{\cup}_{1\leq n\leq\infty}(\operatorname{\mathcal{X}}\times\operatorname{\mathcal{Y}})^{n}\to\boldsymbol{\cup}_{m\leq n}(\operatorname{\mathcal{X}}\times\operatorname{\mathcal{Y}})^{m}\times\boldsymbol{\cup}_{\operatorname{\mathbf{i}}\in\operatorname{\mathscr{P}}(n)}\mathscr{M}(\operatorname{\mathbf{i}})$ outputs a compression set $S^{\operatorname{\mathbf{i}}}$ and a message $\mu$. Then, the reconstruction function $\operatorname{\mathscr{R}}:\boldsymbol{\cup}_{m\leq n}(\operatorname{\mathcal{X}}\times\operatorname{\mathcal{Y}})^{m}\times\boldsymbol{\cup}_{\operatorname{\mathbf{i}}\in\operatorname{\mathscr{P}}(n)}\mathscr{M}(\operatorname{\mathbf{i}})\to\Theta$ is defined such that $A(S)=\operatorname{\mathscr{R}}\quantity(S^{\operatorname{\mathbf{i}}},\mu)$. The forthcoming results rely on a probability distribution over the set of sample-compressed predictors $\overline{\Theta}\subseteq\Theta$. For any sample-compressed predictor $\operatorname{\mathscr{R}}(S^{\operatorname{\mathbf{i}}},\mu)$, this data-independent distribution is expressed as $P_{\overline{\Theta}}\quantity(\operatorname{\mathscr{R}}(S^{\operatorname{\mathbf{i}}},\mu)){=}P_{\operatorname{\mathscr{P}}(n)}(\operatorname{\mathbf{i}})P_{\mathscr{M}(\operatorname{\mathbf{i}})}(\mu)$, with $P_{\operatorname{\mathscr{P}}(n)}$ a probability distribution over $\operatorname{\mathscr{P}}(n)$ and $P_{\mathscr{M}(\operatorname{\mathbf{i}})}$ a probability distribution over $\mathscr{M}(\operatorname{\mathbf{i}})$. Following common practices (Marchand and Sokolova, 2005; Bazinet et al., 2025), we consider hereunder $P_{\operatorname{\mathscr{P}}(n)}(\operatorname{\mathbf{i}}){=}{n\choose\operatorname{|\mathbf{i}|}}^{{}_{-1}}\zeta(\operatorname{|\mathbf{i}|})$, with $\zeta(k){=}\tfrac{6}{\pi^{{}_{2}}}(k{+}1)^{-2}$. Also, given a sequence of indices $\operatorname{\mathbf{i}}$, we set $P_{\mathscr{M}(\operatorname{\mathbf{i}})}$ to a uniform distribution over the messages.
@@
We now present a recent generalization guarantee for sample compression (Bazinet et al., 2025), on which we build our theoretical framework. We use the shorthand notation $\operatorname{\widehat{\mathcal{L}}}_{S}^{{\;\;}_{\operatorname{\operatorname{\mathbf{i}}^{{\rm\bf c}}}}}(\cdot)$ to denote the empirical loss on the training sample that do not belong to the compression set: $S^{\operatorname{\operatorname{\mathbf{i}}^{{\rm\bf c}}}}=S\setminus S^{\mathbf{i}}$; it is mandatory to use $\operatorname{\widehat{\mathcal{L}}}_{S}^{{\;\;}_{\operatorname{\operatorname{\mathbf{i}}^{{\rm\bf c}}}}}(\cdot)$ to obtain an unbiased estimate of the loss, as the predictor $\operatorname{\mathscr{R}}(S^{\operatorname{\mathbf{i}}},\mu)$ relies on $S^{\mathbf{i}}$ and the result wouldn’t hold otherwise.
Theorem 2.1 (Bazinet et al., 2025).
[…]